OpenStudy (anonymous):

the speed of a train is 6mph slower thant the speed of train b. train a travels 190 miles in the same tme it takes train b to travel 220 miles.

hero (hero):

You forgot to ask the actual question

OpenStudy (anonymous):

i did, but i typed it wrong and deleted it.

OpenStudy (anonymous):

speed of train a? speed of train b? I keep getting different answers

OpenStudy (anonymous):

what answers did you get?

OpenStudy (anonymous):

speed*hrs = 220 and (speed-6)*hrs = 190 btw

OpenStudy (anonymous):

5 hours have gone by, because the trains are 30 miles apart

OpenStudy (anonymous):

hence 44 and 38

OpenStudy (anonymous):

once i got 42 and 35

OpenStudy (anonymous):

42-35 = 7 difference wups

OpenStudy (anonymous):

but i think i multiplied wrong

OpenStudy (anonymous):

im working it again to check

OpenStudy (ash2326):

Let speed of Train B be x miles/hr Then speed of train A is (x-6) miles/hr it's given that in equal time train A covers 190 miles and train B covers 220 miles so \[ \frac{190}{x-6}=\frac{220}{x}\] or\[190(x)=220(x-6)\] \[190x-220x=-1320\] \[-30x=-1320\] \[x=44\] so spee of train B is 44miles/ hour and train A's speed is 38 miles/hour

OpenStudy (anonymous):

i did. i multiplied wrong and put the wrong number down in my equation. simple mistakes ugh/ thanks